OpenStudy (yueyue):

Find the limit of f(x)=3/x^4-1 as x approaches 1. The answer cannot be undefined.

OpenStudy (wolfe8):

What you would want to do is factorize the denominator such that you can get 2 or more different fractions where one of them will not have 0 as the denominator when you substitute x.

OpenStudy (anonymous):

is f(x)=\[\frac{ 3 }{x ^{4}-1 }\]

OpenStudy (anonymous):

??

OpenStudy (yueyue):

Yes, that's right.

OpenStudy (anonymous):

ohk so factorize it as \[(x^{2}+1)(x-1)(x+1)\]

OpenStudy (anonymous):

but in the numerator you must have a factor of x-1 else f(x) is undefined

OpenStudy (wolfe8):

I think you can use (x^3 -1/x) x

OpenStudy (yueyue):

Well the original problem asks me to find f(g(x)), where f(x) = 3/x-1 and g(x)= x^4, so there's not a factor of x-1 originally.

OpenStudy (jdoe0001):

can you post a quick screenshot of the material?

OpenStudy (anonymous):

it can be noted that if limiting value tends to infinity ans. is still valid i.e at neighborhood of x-->1 f(x) tends to assume very large values.. but its DISCONTINUOS at x=1

OpenStudy (jdoe0001):

hmm... well, yes, that's correct, so if we stick a positive value to \(\bf x^4\) it will spit out a positive one, if we stick negative ones, it will also spit out positive ones the bigger the value, the greater the denominator, the smaller the fraction if the fraction gets smaller, then that means is approaching 0 that's is true for both one-sided limits though as hantenks correctly said, at 1 is undefined, but doesn't matter, the limit is just an approachment point

OpenStudy (yueyue):

Alright, so in summary my answer is undefined no matter what?

OpenStudy (jdoe0001):

hmm, no, is 0

OpenStudy (debbieg):

Sorry... I don't understand, why do we want to factor \(x^4-1\)? what is the purpose of that? The limit as x->1+ is infinity, and limit as x->1- is -infinity.

OpenStudy (jdoe0001):

hmmm shoot I get something else from the graph :(

OpenStudy (debbieg):

There is a vertical asmptote at x=1. Since the den'r is non-zero, it's just a straightforward one-sided limit, isn't it? (maybe I'm missing something.. won't be the first time...lol)

OpenStudy (anonymous):

i just referred that if there was a factor of x-1 in numerator we should have factorized it

OpenStudy (anonymous):

it=x^4 -1 @DebbieG

OpenStudy (debbieg):

I bet you entered 3/x^4-1 for the graph,, not 3/(x^4-1)... maybe?

OpenStudy (yueyue):

Ok sorry but I'm still not clear on the answer and how to get there.

OpenStudy (jdoe0001):

ohh, I see my mistake.... I was... anyhow, yes.. hhehe the graph is fine, it was just that I was using bigger integers for "x", thus moving away from 1

OpenStudy (debbieg):

ok, wait... isn't the function we're talking about: \[\Large f(x)=\frac{ 3 }{ x^4-1 }\] Or no??

OpenStudy (anonymous):

yes

OpenStudy (yueyue):

Yes

OpenStudy (debbieg):

OK, limits at x=1 differ from the left and from the right. Go to +inf from the right, -inf from the left. Hence, limit does not exist, but the 1-sided limits to, but are infinite. Right?

OpenStudy (jdoe0001):

so yeah,.... if I stick rationals on \(\bf x^4\) it makes the fracton bigger, thus a bigger number off off to \(\bf +\infty \ and \ -\infty\)

OpenStudy (wolfe8):

Here is my solution although it's kinda radical:\[(x + 1)^{2}(x - 1)^{2}-x ^{3}+x ^{2}+2x\] as the denominator. Then substituting x = 1 in the equation will have a definite answer. I haven't checked if I expanded the brackets correctly though.

OpenStudy (debbieg):